OpenStudy (anonymous):

a bank teller cashed a check for $200 using $5 bills then than $10 bills. The bank teller handed 25 more $5 bills than $10 to bills to the customer. Determine how many bills of each domination were handed to the customer.

OpenStudy (anonymous):

f = # of five dollar bills t = # of ten dollar bills f = t + 25 5f + 10t = 200 substitute what f = into the second equation. 5(t + 25) + 10t = 200 5t + 125 + 10t = 200 15t + 125 = 200 15t = 75 t = 5 f = t + 25 so f = 5 + 25 so f = 30 5 ten dollar bills 30 five dollar bills

OpenStudy (anonymous):

thank again Blexting.